# Order Is All You Need for Categorical Data Clustering

### v1

@@ 4. Experiments
Table 2. Statistics of the 12 datasets. $d^{\text{\{nom\}}}$, $d^{\text{\{ord\}}}$, and $n$ are the numbers of nominal attributes, ordinal attributes, and samples, respectively. $k^{*}$ is the number of true clusters.
@@
1 Soybean (small) SB 35 0 47 4
2 Nursery School NS 1 7 12960 4
3 Amphibians AP 6 6 189 2
4 Inflammations Diagnosis DS 0 5 120 2
5 Caesarian Section CS 1 3 80 2
6 Hayes-Roth HR 2 2 132 3
7 Zoo ZO 15 0 101 7
8 Breast Cancer BC 5 4 286 2
9 Lymphography LG 15 3 148 4
10 Tic-Tac-Toe TT 9 0 958 2
11 Australia Credit AC 0 8 690 2
12 Congressional Voting VT 16 0 435 2
4.1. Experimental Settings
Nine counterparts have been selected for the comparative study. The counterparts include traditional K-MoDes (KMD) algorithm (Huang, 1998), popular entropy-based metric LSM (Lin, 1998), and context-based distance measurement CBDM (Ienco et al., 2009). The latter two are representative measures proposed for categorical data, both of which can be embedded into KMD to conduct clustering. Besides, some state-of-the-art methods have also been taken into account, including Jia’s Distance Metric (JDM) (Jia et al., 2016), Unified Distance Metric based Clustering (UDMC) (Zhang and Cheung, 2022b), HD based Clustering (HDC) (Zhang and Cheung, 2022a), projection based clustering Het2Hom (H2H) (Zhang et al., 2022), Distance Learning based Clustering (DLC) algorithm (Zhang and Cheung, 2020), and graph metric based clustering algorithm ADC (Zhang and Cheung, 2023). Please note that we have set the parameters (if any) of these methods to the values recommended in the corresponding paper. In the experiments, each clustering result was recorded based on the average performance of 10 runs.
Three validity indices, i.e., clustering accuracy (CA) (He et al., 2005), ARI coefficients (Rand, 1971)(Gates and Ahn, 2017), and standardized mutual information NMI (Estévez et al., 2009), have been selected to evaluate the clustering performance. Usually, larger values of these indices indicate better clustering performance. We also conduct the Bonferroni Dunn (BD) test [^3] The Friedman test (Friedman, 1937) is performed first to ensure a meaningful BD post-hoc test. (Demšar, 2006) on the performance of different methods and calculate the critical difference (CD) interval to statistically analyze the superiority of our OCL.
[…]
Table 3. Clustering performance evaluated by CA. “$\overline{\textit{AR}}$” row reports the average performance rankings.
Datasets KMD LSM JDM CBDM UDMC DLC H2H HDC ADC OCL
SB 0.8468$\pm$0.15 0.8106$\pm$0.16 0.7894$\pm$0.14 0.7957$\pm$0.15 0.7915$\pm$0.18 0.8000$\pm$0.14 0.9234$\pm$0.12 0.7915$\pm$0.15 0.7660$\pm$0.13 0.9830$\pm$0.05
NS 0.3208$\pm$0.05 0.2975$\pm$0.04 0.2980$\pm$0.04 - 0.3074$\pm$0.05 0.3386$\pm$0.07 0.3113$\pm$0.05 0.3208$\pm$0.05 0.3208$\pm$0.05 0.3573$\pm$0.05
AP 0.5392$\pm$0.03 0.5656$\pm$0.04 0.6159$\pm$0.04 0.5926$\pm$0.03 0.5698$\pm$0.03 0.6032$\pm$0.00 0.5603$\pm$0.03 0.5608$\pm$0.04 0.5529$\pm$0.04 0.6222$\pm$0.00
DS 0.6525$\pm$0.08 0.6525$\pm$0.08 0.7208$\pm$0.12 0.7208$\pm$0.12 0.6867$\pm$0.12 0.6650$\pm$0.12 0.6175$\pm$0.04 0.6867$\pm$0.12 0.6950$\pm$0.12 0.7458$\pm$0.08
CS 0.5462$\pm$0.03 0.5488$\pm$0.06 0.5788$\pm$0.07 0.5875$\pm$0.04 0.5825$\pm$0.05 0.6350$\pm$0.03 0.6075$\pm$0.02 0.5837$\pm$0.05 0.5525$\pm$0.04 0.6313$\pm$0.04
HR 0.3758$\pm$0.02 0.3886$\pm$0.03 0.3917$\pm$0.02 0.3924$\pm$0.04 0.3773$\pm$0.02 0.3697$\pm$0.04 0.3333$\pm$0.00 0.4000$\pm$0.03 0.3758$\pm$0.02 0.4326$\pm$0.02
ZO 0.6624$\pm$0.12 0.6733$\pm$0.11 0.6743$\pm$0.11 0.6752$\pm$0.10 0.6693$\pm$0.10 0.6802$\pm$0.08 0.7059$\pm$0.05 0.6733$\pm$0.10 0.6743$\pm$0.11 0.7792$\pm$0.10
BC 0.5472$\pm$0.04 0.5594$\pm$0.07 0.6392$\pm$0.09 0.6346$\pm$0.09 0.5413$\pm$0.06 0.6140$\pm$0.09 0.5266$\pm$0.06 0.6052$\pm$0.10 0.6185$\pm$0.10 0.6650$\pm$0.05
LG 0.4277$\pm$0.06 0.4446$\pm$0.04 0.4547$\pm$0.06 0.4905$\pm$0.06 0.4466$\pm$0.05 0.5095$\pm$0.08 0.4959$\pm$0.05 0.4959$\pm$0.05 0.5101$\pm$0.04 0.5426$\pm$0.02
TT 0.5569$\pm$0.05 0.5593$\pm$0.04 0.5563$\pm$0.04 0.5408$\pm$0.03 0.5386$\pm$0.03 0.5637$\pm$0.03 0.5640$\pm$0.04 0.5465$\pm$0.04 0.5471$\pm$0.04 0.5785$\pm$0.03
AC 0.7281$\pm$0.12 0.7783$\pm$0.08 0.6335$\pm$0.12 0.6354$\pm$0.12 0.7683$\pm$0.08 0.7488$\pm$0.14 0.7703$\pm$0.08 0.6762$\pm$0.14 0.6293$\pm$0.14 0.8206$\pm$0.08
VT 0.8625$\pm$0.01 0.8699$\pm$0.00 0.8699$\pm$0.00 0.8761$\pm$0.00 0.8667$\pm$0.00 0.8531$\pm$0.09 0.8736$\pm$0.00 0.8736$\pm$0.00 0.8713$\pm$0.00 0.8943$\pm$0.00
$\overline{\textit{AR}}$ 7.58 6.46 5.71 5.13 7.08 4.67 5.33 5.63 6.33 1.08
Table 4. Clustering performance evaluated by ARI (with range [-1,1]). “$\overline{\textit{AR}}$” row reports the average performance rankings.
Datasets KMD LSM JDM CBDM UDMC DLC H2H HDC ADC OCL
SB 0.7638$\pm$0.18 0.7555$\pm$0.19 0.7232$\pm$0.19 0.7407$\pm$0.19 0.7307$\pm$0.22 0.7148$\pm$0.20 0.8650$\pm$0.22 0.7350$\pm$0.19 0.6968$\pm$0.17 0.9620$\pm$0.12
NS 0.0376$\pm$0.03 0.0341$\pm$0.02 0.0283$\pm$0.02 - 0.0393$\pm$0.03 0.0779$\pm$0.08 0.0309$\pm$0.03 0.0376$\pm$0.03 0.0376$\pm$0.03 0.1015$\pm$0.06
AP 0.0022$\pm$0.01 0.0171$\pm$0.02 0.0530$\pm$0.03 0.0322$\pm$0.02 0.0178$\pm$0.02 0.0355$\pm$0.00 0.0076$\pm$0.02 0.0150$\pm$0.02 0.0066$\pm$0.02 0.0524$\pm$0.00
DS 0.1077$\pm$0.13 0.1077$\pm$0.13 0.2347$\pm$0.20 0.2347$\pm$0.20 0.1844$\pm$0.20 0.1556$\pm$0.19 0.0512$\pm$0.05 0.1844$\pm$0.20 0.1921$\pm$0.19 0.2590$\pm$0.16
CS -0.0023$\pm$0.01 0.0088$\pm$0.04 0.0295$\pm$0.04 0.0250$\pm$0.03 0.0248$\pm$0.04 0.0647$\pm$0.03 0.0342$\pm$0.02 0.0234$\pm$0.03 0.0040$\pm$0.02 0.0627$\pm$0.03
HR -0.0081$\pm$0.01 -0.0008$\pm$0.01 -0.0013$\pm$0.01 -0.0010$\pm$0.01 -0.0046$\pm$0.01 -0.0033$\pm$0.01 -0.0149$\pm$0.00 0.0038$\pm$0.02 -0.0064$\pm$0.01 0.0368$\pm$0.01
ZO 0.5787$\pm$0.17 0.6085$\pm$0.17 0.5996$\pm$0.17 0.6091$\pm$0.15 0.6063$\pm$0.16 0.6094$\pm$0.12 0.6442$\pm$0.07 0.6138$\pm$0.15 0.6143$\pm$0.15 0.7536$\pm$0.11
BC 0.0046$\pm$0.02 0.0240$\pm$0.05 0.0811$\pm$0.07 0.0777$\pm$0.07 0.0128$\pm$0.04 0.0523$\pm$0.05 -0.0033$\pm$0.00 0.0605$\pm$0.08 0.0692$\pm$0.08 0.0799$\pm$0.03
LG 0.0843$\pm$0.05 0.0951$\pm$0.03 0.1027$\pm$0.06 0.1432$\pm$0.07 0.1018$\pm$0.04 0.1544$\pm$0.09 0.1715$\pm$0.04 0.1095$\pm$0.06 0.1262$\pm$0.05 0.1552$\pm$0.03
TT 0.0170$\pm$0.03 0.0184$\pm$0.02 0.0181$\pm$0.02 0.0095$\pm$0.01 0.0091$\pm$0.01 0.0156$\pm$0.01 0.0198$\pm$0.02 0.0120$\pm$0.02 0.0124$\pm$0.02 0.0226$\pm$0.02
AC 0.2596$\pm$0.18 0.3336$\pm$0.12 0.1210$\pm$0.15 0.1199$\pm$0.13 0.3092$\pm$0.11 0.3184$\pm$0.22 0.3107$\pm$0.11 0.1913$\pm$0.18 0.1349$\pm$0.18 0.4313$\pm$0.16
VT 0.5247$\pm$0.02 0.5463$\pm$0.01 0.5463$\pm$0.01 0.5648$\pm$0.01 0.5368$\pm$0.01 0.5200$\pm$0.18 0.5572$\pm$0.00 0.5572$\pm$0.00 0.5503$\pm$0.00 0.6207$\pm$0.00
$\overline{\textit{AR}}$ 7.88 5.92 5.50 5.38 6.71 5.08 5.29 5.50 6.42 1.33
4.2. Clustering Performance Evaluation
In this section, we investigate the clustering performance of different algorithms and statistically analyze the superiority of OCL.
4.2.1. Clustering performance on different datasets
The clustering results obtained by different algorithms on various datasets in terms of different validity indices are shown in Tables 3 and 4. The best and second-best results on each dataset are highlighted in bold and underlined, respectively. The observations include the following four aspects: (1) Overall, OCL reaches the best result on most datasets, indicating its superiority in clustering. (2) On the CA index, although OCL does not perform significantly better than the second-best method on the AP and TT datasets, the second-best methods are different on these datasets, but OCL has stable performance on them. (3) On the CA index, although OCL does not perform the best on CS dataset, its performance is close to the best result (the difference is less than 0.005), which still proves OCL’s competitiveness. (4) In terms of the ARI index, there are four datasets on which OCL does not perform the best, namely AP, CS, BC, and LG. However, similarly, the differences between OCL and the best-performing algorithms on AP, CS, and BC datasets are very small. In addition, the result of CBDM is not reported on the NS dataset because all the NS’s attributes are independent of each other, making the inter-attribute dependence-based CBDM fail in measuring distances.
4.2.2. Significance study
We first implement the Friedman (Friedman, 1937) test on the average ranks reported in the last rows in Tables 3 and 4. The corresponding p-values are 0.000017 and 0.000091 respectively, both passing the test under the confidence of 99% (p-value $=0.01$). On this basis, results of the two-tailed BD test (Demšar, 2006) with confidence intervals 0.95 ($\alpha=0.05$) and 0.9 ($\alpha=0.1$) are shown in Figure 2, with the corresponding CD intervals CD_95%=3.4275 and CD_90%=3.1383, respectively. According to the criterion presented in (Demšar, 2006), OCL is considered to perform significantly better than all the counterparts that fall outside the right boundary of the CD intervals. It can be seen that the proposed OCL is significantly superior to all counterparts.
Figure 2. Results of the two-tailed BD tests w.r.t. the CA (a) and ARI (b) performance in Tables 3 and  4, respectively.
4.3. Ablation Studies
To demonstrate the effectiveness of the core components of OCL and to verify our hypothesis on the orders of categorical data attributes, we design several variants of OCL for ablation studies from both algorithm and data perspectives.
From the algorithm aspect, to evaluate the effectiveness of the proposed Order Distance (OD), we compare OCL with its variant OCLI, which uses normalized equidistant order distance without considering the probability distribution of possible values. To evaluate the effectiveness of the Order Learning (OL) mechanism, we further remove the process of order learning from OCLI and only update the order once in the first iteration, which forms another variant OCLII. To completely discard Order Information (OI), we further make OCLII use only the traditional Hamming distance, thus forming the third variant OCLIII. Table 5 compares the clustering performance of OCL and the three variants.
It can be observed that the performance of OCL is superior to all variants. More specifically, OCL outperforms OCLI significantly on 11 out of all 12 datasets. This indicates that OD can effectively utilize the distribution information of the attribute values within different clusters for reasonable order distance measurement. Furthermore, the performance of OCLI is also significantly better than OCLII on all 12 datasets, which verifies the effectiveness of the OL mechanism in searching optimal orders during clustering. In addition, compared to the OCLIII variant that completely disregards OI, OCLII has better performance on 11 out of all 12 datasets, which proves that even a roughly estimated initial order can still significantly and extensively improve clustering performance.
Table 5. CA performance of ablated OCL variants formed by successively ablating the three core components, i.e., Order Distance (OD), Order Learning (OL), and Order Information (OI). The green arrows indicate the performance improvements of an OCL variant over its previous more basic variant. “$\overline{\textit{AR}}$” column reports the average performance rankings.
OCL Components Variants Datasets $\overline{\textit{AR}}$
OD OL OI SB NS AP DS CS HR ZO BC LG TT AC VT
$\checkmark$ $\checkmark$ OCLI   0.9149$\uparrow$   0.3364$\uparrow$   0.6085$\uparrow$   0.7283$\uparrow$   0.6500$\uparrow$   0.3765$\uparrow$   0.7525$\uparrow$   0.5643$\uparrow$   0.4872$\uparrow$   0.5784$\uparrow$   0.7546$\uparrow$   0.8942$\uparrow$   2.00$\uparrow$
$\checkmark$ OCLII   0.8681$\uparrow$   0.3359$\uparrow$   0.6069$\uparrow$ 0.  7125$\uparrow$   0.6388$\uparrow$ 0.3712   0.7000$\uparrow$   0.5636$\uparrow$   0.4453$\uparrow$   0.5762$\uparrow$   0.7545$\uparrow$   0.8828$\uparrow$   2.92$\uparrow$
OCLIII 0.8468 0.3208 0.5392 0.6525 0.5463 0.3758 0.6624 0.5472 0.4277 0.5569 0.7281 0.8625 3.92
From the data aspect, to validate our hypothesis that correct order information is crucial to categorical data clustering, we conduct ablation experiments on datasets containing both nominal and ordinal attributes. We compare the proposed OCL, which learns orders for both the nominal and ordinal attributes, to two different variants. The first one is LNRO (Learn Nominal Reserve Ordinal), which retains the semantic order information of the original ordinal attribute and only learns orders for the nominal attributes. The other one is RNRO (Reserve Nominal Reserve Ordinal), which discards the order learning process for both nominal and ordinal attributes. Table 6 compares the clustering performance of the three variants. It can be observed that on five out of the six datasets, the performance of LNRO is superior to RNRO. This verifies our hypothesis that the nominal attributes also contain potential order information that is effective for clustering. Significantly, we can find that the OCL method outperforms LNRO on five out of the six datasets. This indicates that the orders learned by the proposed mechanism for ordinal attributes are more conducive to clustering than the original semantic orders.
Table 6. CA performance of OCL variants treating Nominal Attributes (NA) and Ordinal Attributes (OA) in different ways.
Learn Order for Variants Datasets $\overline{\textit{AR}}$
OA NA NS AP CS HR BC LG
RNRO 0.3308 0.5984 0.6225 0.3902 0.5276 0.4493 2.83
4.4. Convergence and Efficiency Evaluation
Figure 3. Convergence curve of OCL on all 12 datasets.
Furthermore, to evaluate the execution efficiency of OCL, we construct large synthetic datasets with different numbers of samples and attributes to investigate the execution time of OCL in comparison with other counterparts. The results reported in Figure 4 show that the execution time of OCL is better than UDMC and DLC on large-scale datasets, and better than H2H, UDMC, and CBDM on high-dimensional datasets. Additionally, as the number of samples or attributes increases, the execution time of OCL increases linearly, which is consistent with the time complexity analysis in Section 3.5. In summary, OCL is efficient compared to the state-of-the-art methods and does not incur too much additional computational cost compared to the simplest methods.
Figure 4. Execution time on synthetic datasets with different numbers of samples $n$ and attributes $d$.
4.5. Visualization of Clustering Effect
To demonstrate the cluster discrimination ability of the OCL method, we utilize the distance between attribute values learned by OCL, CBDM, and ADC respectively to encode the attributes of the VT dataset. Subsequently, the t-SNE (Maaten and Hinton, 2008) is introduced to process the encoded data into two-dimensional space and visualize it by marking the samples with real labels in different colors. It can be seen from Figure 5 that OCL has significantly better cluster discrimination ability than CBDM and ADC.
(a) CBDM @@@@ VT
(b) ADC @@@@ VT
(c) OCL @@@@ VT
Figure 5. t-SNE visualization of the VT dataset represented by CBDM, ADC, and OCL.
5. Concluding Remarks
This paper introduces a new finding that knowing the correct order relation of attribute values is critical to cluster analysis of categorical data. Accordingly, we propose the OCL paradigm that allows joint learning of data partitions and order relation of attribute values for accurate clustering. The proposed learning mechanism can infer a more appropriate order among categorical attribute values based on the current clusters, thus yielding more compact clusters based on the updated order. The data partition and order learning are iteratively executed to approach the clustering objective, i.e., maximizing the overall similarity of intra-cluster samples. As a result, more accurate clustering results and insightful orders can be obtained simultaneously. OCL converges quickly with guarantee, and extensive experiments conducted on 12 real benchmark datasets fully demonstrate its superiority in comparison with state-of-the-art clustering methods. A case study has also been provided to show the rationality of the learned order in understanding unlabeled categorical data.
The proposed OCL method is also not exempt from limitations. Although it demonstrates the potential to enhance the performance of mixed data clustering, the deep connection between the treatment of categorical and numerical attributes remains to be explored. In addition, it would also be promising to extend OCL to more complex datasets with non-static distribution, an unknown number of clusters, and extremely imbalanced clusters.

### v3

@@ A.11 Clustering Performance on Numerical Datasets
Table 14: Dataset statistics of the six larger-scale datasets. $d^{\text{\{nom\}}}$, $d^{\text{\{ord\}}}$, $n$, and $k^{*}$ are the numbers of nominal attributes, ordinal attributes, samples, and true clusters, respectively. The full name of the “Coupon” dataset is “In-Vehicle Coupon Recommendation”.
@@
1 Obesity Levels OB 6 2 2111 7
2 Auction Verifications AV 0 6 2043 2
3 Chess CC 36 0 3195 2
4 Coupon CR 7 4 12684 2
5 Bank Marking BM 6 3 45211 2
6 Connect-4 CT4 42 0 67557 3
Table 15: Clustering performance evaluated by CA, NMI, and ARI on the six larger-scale datasets. The best and second-best results on each dataset are highlighted in bold and underlined, respectively. The “$\overline{\textit{AR}}$” row reports the average performance ranks of each method in all 18 comparisons.
Data Metric KMD LSM JDM CBDM UDMC DLC HARR HDC ADC MCDC AMPHM OCL
CT4 CA 0.4000$\pm$0.02 0.3609$\pm$0.01 0.3696$\pm$0.02 0.3651$\pm$0.02 0.3819$\pm$0.03 0.3618$\pm$0.01 0.3924$\pm$0.03 0.3879$\pm$0.03 0.3914$\pm$0.03 0.3514$\pm$0.01 0.3824$\pm$0.00 0.4121$\pm$0.02
ARI -0.0027$\pm$0.00 0.0010$\pm$0.00 0.0001$\pm$0.00 0.0004$\pm$0.00 0.0013$\pm$0.00 0.0012$\pm$0.00 $0.0$$0.0$0005$\pm$0.00 0.0002$\pm$0.00 0.0005$\pm$0.00 0.0150$\pm$0.01
NMI 0.0028$\pm$0.00 0.0029$\pm$0.00 0.0017$\pm$0.00 0.0015$\pm$0.00 0.0036$\pm$0.00 0.0013$\pm$0.00 0.0011$\pm$0.00 0.0016$\pm$0.00 0.0016$\pm$0.00 0.0053$\pm$0.00 0.0023$\pm$0.00 0.0128$\pm$0.01
OB CA 0.3715$\pm$0.03 0.3697$\pm$0.02 0.3594$\pm$0.03 0.3673$\pm$0.03 0.3720$\pm$0.02 0.3724$\pm$0.04 0.3797$\pm$0.02 0.3617$\pm$0.04 0.3765$\pm$0.03 0.3717$\pm$0.02 0.3512$\pm$0.00 0.3935$\pm$0.01
ARI 0.1527$\pm$0.03 0.1588$\pm$0.02 0.1347$\pm$0.04 0.1522$\pm$0.04 0.1580$\pm$0.02 0.1548$\pm$0.04 0.1727$\pm$0.02 0.1417$\pm$0.05 0.1472$\pm$0.04 0.1527$\pm$0.04 0.1672$\pm$0.00 0.1696$\pm$0.01
NMI 0.2256$\pm$0.03 0.2330$\pm$0.02 0.2081$\pm$0.04 0.2262$\pm$0.03 0.2342$\pm$0.02 0.2391$\pm$0.03 0.2644$\pm$0.02 0.2261$\pm$0.04 0.2233$\pm$0.03 0.2466$\pm$0.03 0.2333$\pm$0.00 0.2633$\pm$0.01
AV CA 0.6251$\pm$0.08 0.6035$\pm$0.10 0.6291$\pm$0.12 0.6695$\pm$0.10 0.6065$\pm$0.10 0.6150$\pm$0.06 0.6441$\pm$0.14 0.6364$\pm$0.10 0.6325$\pm$0.12 0.6114$\pm$0.10 0.6125$\pm$0.00 0.6637$\pm$0.07
ARI 0.0073$\pm$0.02 0.0111$\pm$0.02 0.0136$\pm$0.03 0.0365$\pm$0.02 0.0123$\pm$0.03 0.0262$\pm$0.03 0.0214$\pm$0.03 0.0199$\pm$0.02 0.0173$\pm$0.02 0.0209$\pm$0.02 0.0166$\pm$0.00 0.0393$\pm$0.02
NMI 0.0021$\pm$0.00 0.0024$\pm$0.00 0.0024$\pm$0.00 0.0099$\pm$0.01 0.0027$\pm$0.00 0.0093$\pm$0.01 0.0075$\pm$0.01 0.0035$\pm$0.00 0.0034$\pm$0.00 0.0025$\pm$0.01 0.0044$\pm$0.00 0.0110$\pm$0.01
BM CA 0.5730$\pm$0.06 0.5792$\pm$0.04 0.5959$\pm$0.07 0.6011$\pm$0.07 0.5709$\pm$0.03 0.5944$\pm$0.05 0.6064$\pm$0.12 0.5457$\pm$0.03 0.5543$\pm$0.04 0.5573$\pm$0.02 0.5039$\pm$0.00 0.6127$\pm$0.00
ARI 0.0146$\pm$0.03 0.0107$\pm$0.02 0.0183$\pm$0.04 0.0054$\pm$0.05 0.0074$\pm$0.02 0.0178$\pm$0.02 0.0186$\pm$0.01 -0.0037$\pm$0.02 0.0035$\pm$0.02 -0.0007$\pm$0.02 0.0072$\pm$0.00 0.0233$\pm$0.00
NMI 0.0211$\pm$0.00 0.0143$\pm$0.01 0.0135$\pm$0.01 0.0198$\pm$0.01 0.0179$\pm$0.01 0.0121$\pm$0.01 0.0155$\pm$0.01 0.0192$\pm$0.01 0.0204$\pm$0.01 0.0103$\pm$0.01 0.0177$\pm$0.00 0.0194$\pm$0.00
CC CA 0.5562$\pm$0.04 0.5572$\pm$0.03 0.5632$\pm$0.03 0.5433$\pm$0.03 0.5587$\pm$0.02 0.5342$\pm$0.04 0.5248$\pm$0.02 0.5438$\pm$0.03 0.5395$\pm$0.04 0.5668$\pm$0.03 0.5012$\pm$0.00 0.5726$\pm$0.05
ARI 0.0165$\pm$0.01 0.0166$\pm$0.01 0.0189$\pm$0.02 0.0103$\pm$0.01 0.0150$\pm$0.01 0.0099$\pm$0.02 0.0019$\pm$0.01 0.0108$\pm$0.01 0.0113$\pm$0.02 0.0177$\pm$0.01 0.0106$\pm$0.00 0.0285$\pm$0.02
NMI 0.0150$\pm$0.01 0.0149$\pm$0.01 0.0149$\pm$0.01 0.0084$\pm$0.01 0.0119$\pm$0.01 0.0076$\pm$0.01 0.0023$\pm$0.01 0.0148$\pm$0.01 0.0141$\pm$0.02 0.0156$\pm$0.01 0.0144$\pm$0.00 0.0213$\pm$0.02
CR CA 0.5209$\pm$0.02 0.5247$\pm$0.02 0.5185$\pm$0.02 0.5297$\pm$0.02 0.5245$\pm$0.01 0.5488$\pm$0.03 0.5541$\pm$0.00 0.5466$\pm$0.02 0.5504$\pm$0.02 0.5246$\pm$0.02 0.5510$\pm$0.00 0.5590$\pm$0.02
ARI 0.0021$\pm$0.00 0.0033$\pm$0.00 0.0024$\pm$0.00 0.0039$\pm$0.00 0.0030$\pm$0.00 0.0120$\pm$0.01 0.0114$\pm$0.00 0.0101$\pm$0.01 0.0106$\pm$0.01 0.0096$\pm$0.00 0.0122$\pm$0.00 0.0145$\pm$0.01
NMI 0.0017$\pm$0.00 0.0024$\pm$0.00 0.0020$\pm$0.00 0.0022$\pm$0.00 0.0022$\pm$0.00 0.0071$\pm$0.01 0.0115$\pm$0.00 0.0065$\pm$0.00 0.0057$\pm$0.00 0.0066$\pm$0.00 0.0048$\pm$0.00 0.0083$\pm$0.01
$\overline{\textit{AR}}$ 7.58 7.28 8.06 6.92 6.97 6.39 4.94 7.53 7.08 6.75 7.11 1.39
A.12 Additional Large-scale Datasets Results
Six additional larger datasets in different fields have been chosen from the UCI machine learning repository (see http://archive.ics.uci.edu/ml) to further evaluate the clustering performance of the proposed OCL method. The statistics of the newly added datasets are summarized in Table 14. All datasets are pre-processed following the same settings in the submitted paper (see Section 4.1 and Appendix A.3).
Clustering Performance
The clustering performance of different methods is compared on the six additional datasets in Table 15. Consistent with the results in our paper, the proposed OCL outperforms the other methods in general. More detailed observations are provided below: (1) Average performance ranking of OCL is around 1.4 in 18 comparisons formed on 6 datasets under 3 metrics, while the most competitive rivals, i.e., DLC and HARR, rank around 6.4 and 4.9, respectively. The huge ranking gap (larger than 3) intuitively demonstrates the superiority of OCL. (2) All the experimental datasets are real public datasets from various fields, such as healthcare, gaming, recommendation systems, finance, etc. This configuration is sufficient to verify the superiority of OCL without bias. (3) The proposed OCL always ranks first or second in all comparisons (except for NMI performance on the BM dataset) and performs the best in 13 out of 18 comparisons, indicating that the advantage of OCL is very stable and extensive.